OpenStudy (anonymous):

Please help :( How would i do this?? One function, f(x), with two real rational solutions. One function, g(x), with two real irrational solutions. One function, h(x), with two complex solutions. Just the steps, or how i would do this, please. :) Please - Thank you! ♥

OpenStudy (anonymous):

They have to be in standard form too. But, it is ok if they are not. I can do that myself. :)

OpenStudy (ash2326):

@♥Love.Missy24♥ Have you studied quadratic equations?

OpenStudy (anonymous):

Yea - but i don't recall all of it. Plus i have never been good at math. I only passed by like a few points.. :/

OpenStudy (ash2326):

no problem, you can always learn it. A quadratic equation is of the form \[y=ax^2+bx+c\] Solution of quadratic equation is the values of x for which y=0 and this is also a function, we can replace y by f(x), just a denotion \[f(x)=ax^2+bx+c\] Do you follow till here?

OpenStudy (anonymous):

Yea- i think i got that part! ^.^

OpenStudy (ash2326):

Cool, now solution of a quadratic equation is given by this formula \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] this will give us two values, on using + and other using -

OpenStudy (ash2326):

*one

OpenStudy (anonymous):

Okay - i think i remember that part - i didn't like it lol :) So, how would this work with making the function with two real rational solutions?

OpenStudy (ash2326):

yes, let's make our equation easy I'll put a=1 b=0 \[f(x)=x^2+c\] Let's put f(x) =0, to find value of x in terms of c, can you try?

OpenStudy (anonymous):

Ok so it would all be equal to 0

OpenStudy (anonymous):

0 = 1x^2 + 0x + c ?

OpenStudy (ash2326):

yes, can you find x in terms of c from this \[x^2+c=0\]

OpenStudy (anonymous):

? I do not think so. What would be the first step in solving that?

OpenStudy (anonymous):

Sorry :/

OpenStudy (ash2326):

subtract c from both sides

OpenStudy (anonymous):

so it would be x^2 = 0 - c

OpenStudy (ash2326):

yes \[x^2=-c\] Now take root both sides

OpenStudy (anonymous):

ok so x = c

OpenStudy (anonymous):

oh my bad c^2

OpenStudy (ash2326):

you have to take square if \[x^2=4\] then\[x=\pm \sqrt 4\] do the same thing here

OpenStudy (anonymous):

oh? so it would be (i do not know how to make it show like that) x = +-sqrt(c) ??

OpenStudy (ash2326):

yes, it's ok But would you have c in the square root or something else?

OpenStudy (anonymous):

In the square root. Am i on the right track - or so far out in left field :/ Sorry if i made this complicated.

OpenStudy (ash2326):

No problem, I mean we had x^2=4 so I wrote x=+- sqrt (4) now we have \[x^2=\underline {- c}\] There should be something different, shouldn't it?

OpenStudy (anonymous):

yea.

OpenStudy (ash2326):

So what would be the value of x

OpenStudy (anonymous):

4 ?

OpenStudy (ash2326):

How do you say 4?

OpenStudy (anonymous):

cause you said x^2 = 4

OpenStudy (anonymous):

the way to do it is by considering the polynomial: \[f(x) = ax^2 + bx + c = 0\] and knowing that the x values/solution are found with the quadratic formula:\[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] we are only going to observe the term:\[\pm \sqrt{b^2 - 4ac}\] two real rational solutions: when b^2 - 4ac > 0 and sqrt(b^2 - 4ac) is rational two real irrational solutions: when b^2 - 4ac > 0 and sqrt(b^2 - 4ac) is irrational two complex solutions: when b^2 - 4ac < 0 [leading to a negative under the root] lmk if you have questions. i'm pretty sure this is how they want you to do it. you'll just need to find a,b and c values which satisfy the conditions

OpenStudy (ash2326):

Even if x^2=4. x would be the root of 4. That was just for explanation \[x^2=-c=> x=?\]

OpenStudy (anonymous):

Thank you guys very much. I got it - kinda - i will ask you guys if i have any more questions. ♥ Thank you so much!

OpenStudy (ash2326):

Did you understand?

OpenStudy (anonymous):

Yea - kinda - but i have my parents yelling at me about this and i just cant think right now :/ I will ask if i have any more questions.

OpenStudy (ash2326):

Aww:( yes, it would take a bit of work to understand this. Come here back when you have a free mind